OpenStudy (anonymous):

A line passes through the points (-10,-4) and (-1,2) . What is the slope of the line that passes through the two points?

OpenStudy (anonymous):

slope would be the gradient, i.e. y2-y1/x2-x1 so it would be 2-(-4) / -1-(-10) = 6/9 = 2/3